# Structure of activity in multiregion recurrent neural networks

### v1

@@ VII Discussion
Figure 10: Tensor diagrams illustrating the extension of multiregion networks to higher-rank communication subspaces with mixture-of-Gaussians loadings. (a) Form of $(1+\partial_{t})S^{\mu\nu}(t)$ in the multiregion network we have studied. (b) Form of $(1+\partial_{t})S^{\mu\nu}_{r}(t)$ in the aforementioned extension (in this case, the currents acquire an additional $r$ index that runs over rank-one components). The diagram has two new internal lines corresponding to contractions over mixture components ($c$) and rank-one components ($r^{\prime}$).
In our model, modifying the state of the network, for example, by changing whether a region is in routing or non-routing mode, requires altering synaptic couplings. This could occur dynamically through neuromodulation. An interesting alternative is to use external inputs to control routing. For instance, consider a region that cannot transmit signals due to disorder-induced chaotic fluctuations of the units. If an external stimulus suppresses these fluctuations, the region could gain the ability to transmit signals. This behavior aligns with the experimental observation that neuronal variability quenches at stimulus onset, and suggests that this reduction facilitates cross-region communication [27]. To model this, we can integrate additional input vectors into our framework. This would expand the mean-field equations to include new effective interactions encoding the overlap between the vectors defining the recurrent connectivity and the input vectors. A similar approach was explored by Barbosa et al. [22], who studied a two-region network with a specific configuration of the effective interactions.
A notable aspect of our model and theoretical approach is its alignment with existing methods for neural-data analysis. Specifically, the technique developed by [9] for analyzing multiregion neural recordings involves training a recurrent network to mimic the data, then decomposing the activity in terms of cross-region currents. Intriguingly, our model’s low-dimensional mean-field dynamics offer a closed description in these currents, rather than relying solely on single-region quantities such as two-point functions. This alignment strongly supports the use of current-based analyses in neural data interpretation. Furthermore, our model could be adapted to fit multiregion neural data using approaches akin to those in Valente et al. [56]. Subsequently reducing the model to the mean-field description we derived could provide insights into the dynamics of the fitted model. This positions our work as a bridge connecting practical recurrent-network-based data analysis methods to a deeper analytical understanding of network dynamics.
Another data-driven application of our framework lies in analyzing large-scale connectome data [57]. For example, large-scale reconstructions of neurons and their connections are now available for flies [58, 59], parts of the mammalian cortex [60], and in some other organisms [61]. For connectome datasets where regions are identified, the cross-region connectivity could be approximated as having a low-rank structure, allowing for a reduction using our mean-field framework. This enables a comparison of predicted neuronal dynamics with recorded activity. In scenarios where regions are not already defined, our framework suggests solving the “inverse problem,” namely, determining a partitioning of neurons into regions such that the cross-region connectivity is well approximated by low-rank matrices. Developing a special clustering algorithm for this purpose and applying it to connectome data, such as from the fly, would be interesting. Even in cases where anatomical knowledge suggests certain region definitions, identifying “unsupervised regions” based on the assumption of low-rank cross-region interactions could offer an interesting new functional perspective on regional delineation.
VIII Acknowledgements
We are extremely grateful to L.F. Abbott for his advice on this work. We thank Albert J. Wakhloo for comments on the manuscript. We acknowledge helpful discussions with Ashok Litwin-Kumar and members of his group.
Appendix A Locating multiregion networks in the space of low-rank mixture-of-Gaussians networks
@@
Here, we demonstrate via an explicit construction that multiregion networks are a special case of the low-rank mixture-of-Gaussians model, a framework extensively studied in the literature [54, 55, 56, 62, 63, 64]. We demonstrate this mapping in the disorder-free case. Consider a low-rank network where the rank-one terms are indexed by $r$ (or $r^{\prime}$), neurons are indexed by $i$ (or $j$), and the coupling matrix $W_{ij}$ is defined as
@@
The components of the vectors $v^{r}_{i}$ and $w^{r}_{i}$ follow a mixture-of-Gaussians distribution with i.i.d. sampling across the neuron index, $i$. Each mixture component has zero mean. The second-order statistics are given by
[…]
where $\left\langle\cdot\right\rangle_{c}$ denotes an average in mixture component $c$. We assume that all mixture components have equal probability. With these definitions, the mean-field equations were shown to be
$(1+\partial_{t})\kappa^{r}(t)$ $=\sum_{r^{\prime}}\left[\sum_{c}\psi^{c}(t)t^{rr^{\prime}}[c]\right]\kappa^{r^{\prime}}(t),$ (42)
@@
where $\psi(\Delta)$ is given for the error-function nonlinearity by Eq. 15b.
Toward making a multiregion network emerge from these equations, we consider $R^{2}$ rank-one terms and $R$ mixture components, and substitute $r\rightarrow(\mu,\nu)$, $r^{\prime}\rightarrow(\rho,\sigma)$, and $c\rightarrow\alpha$. The second-order statistics are constructed as follows:

### v3

@@ A.10 DMFT with inputs
Figure 11: Tensor diagrams illustrating the extension of multiregion networks to higher-rank communication subspaces with mixture-of-Gaussians loadings. (a) Form of $(1+d/dt)S^{\mu\nu}(t)$ in the multiregion network we have studied. (b) Form of $(1+d/dt)S^{\mu\nu}_{r}(t)$ in the aforementioned extension (in this case, the currents acquire an additional $r$ index that runs over rank-one components). The diagram has two new internal lines corresponding to contractions over mixture components ($c$) and rank-one components ($r^{\prime}$).
A.12 Locating multiregion networks in the space of low-rank mixture-of-Gaussians networks
Setting aside disorder, our model involves a blockwise low-rank coupling matrix, an embodiment of a broader idea where neurons are assigned group identities and coupling statistics are based on these identities. Another embodiment of this idea is the low-rank mixture-of-Gaussians model proposed by [35, 48], where the coupling matrix is a sum of rank-one outer products with mixture-of-Gaussians loadings. In this framework, each neuronal group corresponds to a Gaussian mixture component. Our multiregion network model, of $R$ regions, is a special case of a low-rank mixture-of-Gaussians network with rank $R^{2}$ and $R$ mixture components.
@@ A.12 Locating multiregion networks in the space of low-rank mixture-of-Gaussians networks
Here, we demonstrate via an explicit construction that multiregion networks are a special case of the low-rank mixture-of-Gaussians model. Consider a low-rank network where the rank-one terms are indexed by $r$ (or $r^{\prime}$), neurons are indexed by $i$ (or $j$), and the coupling matrix $W_{ij}$ is defined as
@@
The components of the vectors $v^{r}_{i}$ and $w^{r}_{i}$ follow a mixture-of-Gaussians distribution with i.i.d. sampling across the neuron index, $i$. Each mixture component has zero mean. The second-order statistics are defined by
$\left\langle w^{r}_{i}v^{r^{\prime}}_{i}\right\rangle_{c}$ $=Ct^{rr^{\prime}}[c],$ (43)
$\left\langle v^{r}_{i}v^{r^{\prime}}_{i}\right\rangle_{c}$ $=Cu^{rr^{\prime}}[c],$ (44)
where $\left\langle\cdot\right\rangle_{c}$ denotes an average in mixture component $c$ and $C$ is the number of mixture components. We assume that all mixture components have equal probability. With these definitions, the mean-field equations were shown in [35, 48] to be
$\left(1+\frac{d}{dt}\right)\kappa^{r}(t)$ $=\sum_{r^{\prime}}\left[\sum_{c}\psi^{c}(t)t^{rr^{\prime}}[c]\right]\kappa^{r^{\prime}}(t),$ (45)
@@
where $\psi(\Delta)$ is given for the error-function nonlinearity by SI Eq. 23a.
Toward making a multiregion network emerge from these equations, we consider $R^{2}$ rank-one terms and $R$ mixture components, and substitute $r\rightarrow(\mu,\nu)$, $r^{\prime}\rightarrow(\rho,\sigma)$, and $c\rightarrow\alpha$ (the purpose of the last replacement is simply to use a Greek letter for consistency). The second-order statistics are constructed as follows:
@@
Furthermore, as described in the main text, multiregion networks can themselves be generalized to have rank-$K$ communication subspaces and $C$ Gaussian-mixture components. This extension can be captured by a low-rank mixture-of-Gaussians construction with $KR^{2}$ rank-one terms and $CR$ mixture components. Compared to generic networks in the low-rank mixture-of-Gaussians class, multiregion networks possess a far greater degree of structure. An interesting question is whether the inductive bias corresponding to multiregion networks is advantageous in constructing models within this class.